A farmer needs 65 gallons of weed killer solution. He has 35% weed killer solution which is too weak for the job and a 60% weed killer solution which is too strong. How many gallons of each solution should he mix to obtain 70 gallons of a 50% solution?
Answer by stanbon(75887)

A farmer needs 65 gallons of weed killer solution.
He has 35% weed killer solution which is too weak for the job and a 60% weed killer solution which is too strong.
How many gallons of each solution should he mix to obtain 70 gallons of a 50% solution?
----------------------------
active + active = active ingrediant
0.35x + 0.60(70-x) = 0.50*70
-------------------------------------
35x + 60*70 - 60x = 50*70
-25x = -10*70
x = 28 (amt. of 35% killer needed in the mixture)
70-x = 42 ( amt. of 60% killer needed in the mixture)
====================================
Cheers,
Stan H.